OpenStudy (anonymous):

find the slope x+6y=12

hero (hero):

-1/6

hero (hero):

y = -x/6 + 2

OpenStudy (anonymous):

\[6y=12-x\] \[y=-\frac{1}{6}x+2\]

OpenStudy (anonymous):

or use implicit differentiation: 1 + 6y' = 0 y' = -1/6